OpenStudy (anonymous):

I need help with this absolute value problem. I know how to do it but just not how to enter it in IMATHAS if any one else uses that. Or if you could just show me how to plug it in on the website. l3x+3l=13 PLZ AND THX!

OpenStudy (anonymous):

\[\left| 3x+3\right|\]=13

OpenStudy (anonymous):

solve two equations \[3x+3=13\] and \[3x+3=-13\] you get two different answers

OpenStudy (anonymous):

i am not sure why you would need to use anything other than a pencil, but if you are dying to do this one the web, try this http://www.wolframalpha.com/input/?i= |3x%2B3|%3D13

OpenStudy (anonymous):

thanks.......I still can't input it in the system i don't know why.. it's getting me upset lol